Istituto ed Osservatorio Astronomico Other Publications and Internal Reports

1997-1993

(written Jul 2 1997)

[1] F. Damiani, G. Micela & S. Sciortino, “Problems and Solutions in deriving the Log N - Log S of sources detected in P SPC images with a wavelettransform algorithm,” 1997, Paper presented at the workshop “X-ray Surveys”, Potsdam 17-22 Giugno 1997.

[2] S. Sciortino, F. Damiani, G. Micela, F. Favata & J. Pye, “An Extended Galactic Plane Survey of the Third Galactic Quadrant (180 ¡ l ¡ 2 80) using a wavelet-transform detection algorithm,” 1997, Paper presented at the workshop “X-ray Surveys”, Potsdam 17-22 Giugno 1997.

[3] F. Favata, G. Micela & S. Sciortino, “The evolutionary status of highlithium cool dwarfs in the solar neighborhood,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997.

[4] F. Favata, S. Sciortino, F. Damiani & G. Micela, “X-ray emitting PMS stars in upper SCO-CEN: the x-ray faint population,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997.

[5] E. Flaccomio, C. Corbally, F. Damiani, F. Favata, F.R. Harnden Jr., G. Micela, S. Sciortino, J. Schachter & A. Tomaney, “X-ray and optical observations on the star forming region NGC 2264,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997.

[6] G. Micela, F. Favata, R. Pallavicini, J. Pye & S. Sciortino, “X-ray observations of ζ Sculptoris,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997.

[7] F. Morale, G. Massone & M.G. Lattanzi, “Proper motions and photographic photometry in the field of the southern open cluster NGC 2516,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997 .

[8] R. Pallavicini, L. Pasquini & S. Randich, “Lithium in solar-type stars in the old cluster M67,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997 .

[9] G. Peres, F. Reale, S. Orlando, R. Rosner & H. Hudson, “Sun as a star: Yohkoh data analysis for the solar-stellar connection,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997 .

[10] J.P. Pye, J.E. Morley, R.S. Warwick, J. Pilkington, G. Micela, S. Sciortino & F. Favata, “A ROSAT deep sky survey of the galactic plane 180 degree ¡l ¡280 degree,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997 .

[11] F. Reale, G. Micela, G. Peres, R. Betta & S. Serio, “Stellar flaring loops,” 1997, Presented at the workshop ”Cool stars in clusters and associations: magnetic activity and age indicators”, Palermo, 20-23 May 1997.

[12] S. Sciortino, G. Micela, F. Reale, V. Kashyap, R. Rosner & F.R. Harnden, Jr, “An Analysis of X-ray Flares in Pleiades Stars.,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996, 277 , Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[13] G. Micela, S. Sciortino & J. Pye, “Coronal Properties of Nearby Old Disk and Halo dM Stars,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996 , 479, Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[14] F. Reale, G. Peres & H. Hudson, “Hydrodynamic Modeling of Flares Well-Observed by Yohkoh/SXT,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands , 1996, 311, Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[15] R.M. Betta, G.Peres, S.Serio & F.Reale, “A regridding algorithm for high resolution hydrodynamics of flaring coronal loops,” in Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity, Y. Uchida, T. Kosugi & H.S. Hudson, eds., Kluwer Academic Publishers, Dordrecht, The Netherlands , 1996, 473, Presented at the IAU Colloquium 153 - Magnetodynamic Phenomena in the Solar Atmosphere - Prototypes of Stellar Activity - Tokyo (Japan) 22-26 May 1995.

[16] A. Ciaravella, A. Maggio, G. Peres & S. Serio, “Detailed Fitting of Coronal X-ray Spectra with Loop Models: Model Characteristics and Simulations,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series , 1996 , 257, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[17] F. Reale, G. Peres & S. Serio, “Radiatively Cooling Downdrafts as the Origin of Redshifts in Transition Region Lines,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series , 1996 , 149 , Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[18] A. Maggio, G. Peres, J.P. Pye, S.T. Hodgkin & J.E. Morley, “Loop Modeling of ROSAT/PSPC Spectra of Hyades and Field F-Type Stars,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series , 1996 , 269, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[19] M. Barbera, F. Bocchino, F. Damiani, F. Favata, F.R. Harnden, Jr., G. Micela & S. Sciortino, “ROSAT PSPC/HRI Observation of the Young Open CLuster NGC2422,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series, 1996 , 355, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[20] V. Kashyap, R. Rosner, F.R. Harnden, Jr., G. Micela & S. Sciortino, “Constraining the Low-Mass Stellar Mass-Function,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series, 1996 , 365, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[21] F. Damiani, G. Micela & S. Sciortino, “X-ray Variability of T Tauri Stars in the rho Oph Region,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Series, 1996 , 423, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[22] F.R. Harnden, Jr., J.P. Caillault, F. Damiani, V. Kashyap, G. Micela, C. Prosser, R. Rosner, S. Sciortino & J. Stauffer, “ROSAT HRI Survey of the Pleiades,” in Cool stars, stellar systems and the sun - 9th Cambridge Workshop, A.K. Dupree & R. Pallavicini, eds., Astronomical Society of the Pacific Conference Serie, 1996 , 359, Presented at the 9th Cambridge Workshop on Cool Stars, Stellar Systems and the Sun, Florence (Italy), 3-6 October 1995.

[23] R. Mewe, J.S. Kaastra, S.M. White & R. Pallavicini, “The Corona of AB Dor: Temperature and Abundance Determinations with ASCA and EUVE,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K.Dupree, eds., 1996, 273.

[24] A. Ortolani, R. Pallavicini & G. Tagliaferri, “X-ray Spectroscopy of RS CVn Binaries: the EXOSAT and SSS Spectra Revisited,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K. Dupree, eds., 1996, 279.

[25] A. Panarella, G. Tagliaferri & R. Pallavicini, “A Study of a Sample of Cool Stars X-ray Selected with EXOSAT using ROSAT Data,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K. Dupree, eds., 1996 , 623 .

[26] S. Randich, N. Aharpour, R. Pallavicini, C.F. Prosser, J.R. Stauffer & J.H.M.M. Schmitt, “Lithium Abundance in the IC 2602 Open Cluster,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K. Dupree, eds., 1996, 379.

[27] G. Tagliaferri, S. Covino, F. Haardt, T.A. Fleming & R. Pallavicini, “An ASCA Observation of HD 35850,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K. Dupree, eds., 1996, 295.

[28] S.M. White, R. Pallavicini & J. Lim, “The X-ray Spectra of Flares from AB Doradus,” in 9th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, R. Pallavicini & A.K. Dupree, eds., 1996, 299.

[29] F. Damiani, A. Maggio, G. Micela & S. Sciortino, “A Source Detection Method for ROSAT/PSPC X-ray Images based on Wavelet Transforms,” in Proceedings of the meeting ”Astronomical Data Analysis Software and Systems V”, G.H. Jacoby & J. Barnes, eds., 1996, 143, Presented at ”ADASS 95”, Tucson, 22-26 October 1995.

[30] G. Mackie, G. Fabbiano, M. Barbera, F. Bocchino, F. Damiani, A. Maggio, G. Micela, S. Sciortino, F. R. Harnden,Jr., D.-W. Kim & P. Ciliegi, “A Reduction and Analysis Pipeline for ROSAT PSPC Data,” in Proceedings of the meeting ”Astronomical Data Analysis Software and Systems V”, G.H. Jacoby & J. Barnes, eds., 1996, 179, Presented at ”ADASS 95”, Tucson, 22-26 October 1995.

[31] G. Tagliaferri, S. Covino, R. Pallavicini & Y. Uchida, “ASCA Observation of Coronal X-ray Emission from the Dwarf Star HD 35850,” in X-ray imaging and Spectroscopy of Cosmic Hot Plasmas, F. Makino, ed., 1996, in press, Presented at X-ray Imaging and Spectroscopy of Cosmic Hot Plasmas, ASCA 3rd Anniversary, Tokyo, 11-14 March 1996.

[32] F. Reale, G. Peres & R. Betta, “Diagnostics of solar flares from hydrodynamic modeling: using the decay phase to derive the loop dimensions,” in Magnetic Reconnection in the Solar Atmosphere, B. Bentley & J. Mariska, eds., 1996, 232 , Proceedings Yohkoh Conference on Observations of Magnetic Reconnection in the Solar Atmosphere, Bath, 20-22 March 1996.

[33] . Chinnici, La ”Societa’ degli Spettroscopisti Italiani” e la fondazione di ”The Astrophysical Journal” nelle lettere di C.E. Hale a P. Tacchini, 1996, Presented at the XVI Congresso Nazionale di Storia della Fisica e dell’Astronomia, Como, 24-25 maggio 1996.

[34] G. Fodera’ Serio & D. Randazzo, “Astronomi italiani dall’Unita’ d’Italia ai giorni nostri: un primo elenco,” 1996, Presented at the XVI Congresso Nazionale di Storia della Fisica e dell’Astronomia, Como, 24-25 maggio 1996 .

[35] F. Damiani, A. Maggio, G. Micela & S. Sciortino, “Statistical Properties of Wavelet Transforms Applied to X-ray Source Detection,” in Proceedings of the meeting ”Statistical Challenges in Modern Astronomy II”, G.J. Babu & E.D. Feigelson, eds., 1996, in press.

[36] A. Collura, M. Barbera, S. Sciortino & S. Serio, “The XACT facility at the Osservatorio Astronomico di Palermo ”G.S. Vaiana”,” in Proceedings of the workshop on High Throughput X-ray Spectroscopy, 1996, Cambridge (MA, USA), 1996, in press, Presented at the Workshop on High Throughput X-ray Spectroscopy, September 30 - October 1, 1996, Cambridge (MA, USA).

[37] A. Collura, M. Barbera, S. Sciortino & S. Serio, “The EUV/X-ray Astronomy Calibration and Testing Facility at the Osservatorio Astronomico di Palermo ”G.S. Vaiana”,” 1996, Presented at the New Generation X-ray Observatories Workshop, September 1996, Leicester (U.K.).

[38] S.M. White, R. Pallavicini, R. Mewe, J.S. Kaastra & J. Lim, “Abundances and Flares in the ASCA Observation of the Young K0 Star AB Doradus,” in X-ray Imaging and Spectroscopy of Cosmic Hot Plasmas, F. Makino, ed., 1996, in press.

[39] N. Aharpour, R. Pallavicini, L. Pasquini & S. Randich, “High-resolution Optical Spectroscopy of Post-T Tauri Stars,” 1996, Presented at the 4th French-Italian Meeting on ”Young Stellar Objects”, Gif-sur-Yvette, 9-10 September 1996 .

[40] E. Flaccomio, F. Damiani, G. Micela, S. Sciortino, F. Favata, F.R. Harnden,Jr., J. Schachter, A. Tomaney & C. Corbally, “X-ray and Optical Study of NGC 2264,” 1996, Presented at the 4th French-Italian Meeting on ”Young Stellar Objects”, Gif-sur-Yvette, 9-10 September 1996.

[41] M. Barbera, A. Collura & A. Dara, “Synchrotron X-ray transmission measurements of Lexan and Lexan+Aluminum samples at BESSY: preliminary report,” Osservatorio Astronomico di Palermo G.S.Vaiana, OAPA-R1-1996, 12 Febbraio, 1996, 1996.

[42] F. Damiani, “AXAF LETGS order sorting,” Osservatorio Astronomico di Palermo G.S.Vaiana, OAPA-R2-1996, 23 Febbraio, 19 96, 1996.

[43] M. Barbera & A. Collura, “UV transmission measurements on two JET-X test filter samples,” Osservatorio Astronomico di Palermo ”G.S. Vaiana”, OAPA-R7-1996, 30 ottobre 1996, 1996.

[44] M. Barbera, A. Collura & S. Jouteux, “Preliminary results on X-ray shadowgraphs conducted at the XACT facility on test thin and medium filters for EPIC-XMM,” Osservatorio Astronomico di Palermo G.S. Vaiana, OAPA-R5-1996, September 23, 1996, 1996.

[45] M. Barbera, A. Collura & A. Dara, “Preliminary UV transmission measurements of HRC imaging UV/Ion shield HRC-1007-1 Rev B S/N 002,” Osservatorio Astronomico di Palermo G.S. Vaiana, OAPA-R3-1996, March 22, 1996, 1996.

[46] F. Bocchino, “Preliminary results on narrow band imagery and spatially resolved spectrophotometry of selected filaments in the Vela Supernova Remnant,” Osservatorio Astronomico di Palermo G.S. Vaiana, OAPA-R4-1996, September 16, 1996 , 1996.

[47] F. Favata, G. Micela & S. Sciortino, “Lithium in Nearby Main-Sequence Solar-Type Stars,” in The Light Element Abundances, P. Crane, ed., 1995, 268, Presented at the ESO/EIPC Workshop “High Resolution Spectroscopy”, held in Marciana Marina, Isola d’Elba, May 21-26 1994.

[48] M. Barbera, A. Collura, F. Favata, G. Inzerillo & F. Mirabello, “Simple Gas Scintillation Proportional Counter soft X-ray detector for laborat ory usage,” in Atti Accademia Peloritana dei Pericolanti, Classe I di Scienze Fisiche, Matematiche e Naturali #LXXII, 1995, 661 , Presented at the 5 Conferenza Scientifica Triennale del Comitato Regionale di ricerche nucleari e di struttura della materia.

[49] F.R. Harnden, Jr., J.P. Caillault, F. Damiani, V. Kashyap, G. Micela, C. Prosser, R. Rosner, S. Sciortino & J. Stauffer, “HRI Observations of the Pleiades,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zimmermann, J.E. Trumper & H. Yorke, eds., 1995 , 43, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[50] F. Bocchino, M. Barbera & S. Sciortino, “Time Screening Optimization Algorithm for ROSAT PSPC/HRI Observations,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zimmermann, J.E. Trumper & H. Yorke, eds., 1995, 673, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[51] F. Bocchino, A. Maggio & S. Sciortino, “On the Study of NEI Regime in Supernova Remnants: Emission Model and ROSAT PSPC Spectral Fitting Simulations,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zinnermann, J.E. Trumper & H. Yorke, eds., 1995 , 235, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[52] F. Damiani, A. Maggio, G. Micela & S. Sciortino, “A Detection Method for X-ray Images based on Wavelet Transforms: the Case of ROSAT PSPC,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zinnermann, J.E. Trumper & H. Yorke, eds., 1995, 679, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[53] F. Damiani, G. Micela & S. Sciortino, “X-ray Variability of T Tauri Stars in the rho Oph Region,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zinnermann, J.E. Trumper & H. Yorke, eds., 1995 , 27, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[54] A. Collura, M. Barbera, S. Sciortino & . S. Serio, “The EUV/X-ray Astronomy Calibration and Testing Facility at the Osservatorio Astronomico di Palermo ”G.S. Vaiana”,” 1995, Presented at the Jenam Meeting held in Catania (Sicily), 25-29 September 1995.

[55] S.E. Pearce, J.E. Lees, J.F. Pearson, G.W. Fraser, A.N. Brunton, K.A. Flanagan, A. Kenter, M. Barbera, V. Dhanak, A. Robinson & D. Teehan, “Synchrotron calibration of alkali halide coated microchannel plate detectors in the 50-350 and 2000-6000 eV bands,” Proceedings SPIE 2518 (1995), 322, Paper presented at SPIE Conference on ”EUV, X-ray and Gamma-ray Instrumentation in Astronomy VI”, San Diego (California, USA), 9-14 July 1995.

[56] R. Pallavicini & S. Randich, “Lithium in Tidally Locked Binaries,” in The Light Element Abundances, P. Crane, ed., 1995, 311.

[57] S. Randich, R. Pallavicini, L. Pasquini & R. Gratton, “Lithium Abundances in Subgiant Stars,” in The Light Element Abundances, P. Crane, ed., 1995, 284.

[58] S. Randich, R. Pallavicini, L. Pasquini & R. Gratton, “Lithium Abundances in Pop I Subgiants,” Memorie S.A.It. 66 (1995), 383.

[59] S.M. White, R. Pallavicini & J. Lim, “ASCA X-ray Spectra of Quiescent and Flaring Emission from AB Doradus,” in Flares and Flashes, J. Greiner, H. Duerbeck & R. Gershberg, eds., 1995, 168 .

[60] G. Mackie, G. Fabbiano, F. Bocchino, P. Ciliegi, F. Damiani, F. R. Harnden,Jr., D.-W. Kim, A. Maggio, G. Micela & S. Sciortino, “Early-Type Galaxies: Automated Reduction and Analyis of ROSAT PSPC Data,” in RoentgenStrahlung from the Universe (MPE Report 263), H.U. Zimmermann, J.E. Trumper & H. Yorke, eds., 1995, 385, Presented at ”RoentgenStrahlung from the Universe”, Wurzburg (Germany), 25-29 September 1995.

[61] M.V. Zombeck, J.H. Chappell, A. Kenter, R.W. Moore, S.S. Murray, G.W. Fraser & S. Serio, “The High Resolution Camera (HRC) on the Advanced X-ray Astrophysics Facility (AXAF),” in SPIE Proceedings #2518, 1995 , 96.

[62] M. Barbera, A. Collura & A. Dara, “Preliminary soft X-ray transmission measurements of HRC imaging UV/Ion shield HRC-1007-1 Rev B S/N # 002 ,” Osservatorio Astronomico di Palermo G.S.Vaiana, OAPA-R3-1995, 31 Ottobre, 1995, 1995.

[63] F. Favata, M. Barbera, G. Micela & S. Sciortino, “The Study of the Stellar Activity-Age Correlation in Field Star Sample Through Lithium Abundance Studies,” in Frontiers of Space and Ground-based Astronomy, W. Wamsteker, M. S. Longair & Y. Kondo, eds., Kluwer Academic Publisher, 1994, 589.

[64] F. Bocchino, A. Maggio & S. Sciortino, “X-ray Emission from the Vela SNR Shock Region: Spectral Fitting with a Non-Equilibrium Ionization Model,” in The Soft X-ray Cosmos (AIP Conference Proceedings 313), Schlegel E. M. & Petre R., eds., 1994, 309, Proceedings of the ROSAT Science Symposium, Maryland, 4-7 Nov. 1993.

[65] A. Maggio, S. Sciortino & F. R. Harnden Jr., “Simulated PSPC Spectral Fits of Coronal X-ray Sources: Limitation in the Case of Low Counting Statistics,” in The Soft X-ray Cosmos (AIP Conference Proceedings 313), Schlegel E. M. & Petre R., eds., 1994, 244, Proceedings of the ROSAT Science Symposium, Maryland, 8-10 Nov. 1993.

[66] F. Reale, A. Maggio, A. Ciaravella & G. Peres, “ASAP: a Systematic Approach to Plasma Spectral Synthesis,” Space Science Reviews 70 (1994), 211, Proceedings of the Second SOHO Workshop ”Mass Supply and Flow in the Solar Corona”, Marciana Marina, Septembre 27 - October 1, 1993.

[67] S. Orlando, S. Serio & G. Peres, “Stationary Flows in Coronal Loops,” Space Science Reviews 70 (1994), 203, Proceedings of the Second SOHO Workshop ”Mass Supply and Flow in the Solar Corona”, Marciana Marina, Septembre 27 - October 1, 1993.

[68] F. Damiani, G. Micela, S. Sciortino & F. R. Harnden, Jr., “Rosat HRI Observations of T-Tauri Star Pairs in Taurus Auriga,” in The Soft X-ray Cosmos (AIP Conference Proceedings 313), Schlegel E. M. & Petre R., eds., 1994, 269, Proceedings of the ROSAT Science Symposium, Maryland, 8-10 Nov. 1993.

[69] V. Kashyap, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “Evaluation of Source Counts And Upper Limits in Crowded ROSAT PSPC Fields,” in The Soft X-ray Cosmos (AIP Conference Proceedings 313), Schlegel E. M. & Petre R., eds., 1994, 239, Proceedings of the ROSAT Science Symposium, Maryland, 8-10 Nov. 1993.

[70] S. Orlando, S. Serio & G. Peres, “Siphon Flows in Stratified Coronal Loops,” in Solar Coronal Structures, V. Rusin, P. Heinzel & J.-C. Vial, eds., VEDA, 1994, 185, Proceedings of the IAU 144 Colloquium ”Solar Coronal Structures”, held in Tatranska Lomnica, Slovacchia, September 20-24, 1993.

[71] B. Sylwester, J. Sylwester, F. Reale & S. Serio, “Manifestation of Multiple Energy Release on the Evolution of Flares in Diagnostic Diagrams,” in Solar Coronal Structures, V. Rusin, P. Heinzel & J.-C. Vial, eds., VEDA, 1994, 257, Proceedings of the IAU 144 Colloquium ”Solar Coronal Structures”, held in Tatranska Lomnica, Slovacchia, September 20-24, 1993.

[72] F. Reale, G. Peres & S. Serio, “Impulsive Heating of Coronal Loops,” in Solar Coronal Structures, V. Rusin, P. Heinzel & J.-C. Vial, eds., VEDA, 1994, 215, Proceedings of the IAU 144 Colloquium ”Solar Coronal Structures”, held in Tatranska Lomnica, Slovacchia, September 20-24, 1993.

[73] F. R. Harnden, Jr., J. Schachter, G. Micela, S. Sciortino & F. Favata, “HRI observation of an Einstein IPC Slew-survey Source near S Mon O star.,” 1994, Presented at the Meeting of the High Energy Division of the AAS, held in Napa Valley (CA), Nov 1-4, 1994.

[74] R. Pallavicini, “Lithium, Activity and Post-main Evolution,” in 8th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, J.-P. Caillault, ed., 1994, 244.

[75] G. Tagliaferri, G. Cutispoto, R. Pallavicini, L. Pasquini & S. Randich, “Optical Studies of Cool Stars Discovered by EXOSAT: Lithium Abundances and Halpha Spectroscopy,” in 8th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, J.-P. Caillault, ed., 1994, 303.

[76] G. Cutispoto, M. Rodono‘, R. Pallavicini & G. Tagliaferri, “Photometry of Serendipitous X-ray Sources Detected by EXOSAT. II An Update,” in 8th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, J.-P. Caillault, ed., 1994, 687.

[77] L. Pasquini, B. Edvardsson, Q. Liu & R. Pallavicini, “Li abundances in Nearby Solar-Like Stars,” in 8th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, J.-P. Caillault, ed., 1994, 294.

[78] S.M. White, R. Pallavicini & J. Lim, “The Radio Spectra of Weak-Lined T Tauri Stars,” in 8th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, J.-P. Caillault, ed., 1994, 501.

[79] M. Barbera, “Soft X-ray/UV plastic film transmission measurements conducted at the Space Sciences Laboratory, University of California at Berkeley: instrumental configuration, measurement techniques, and preliminary results,” Smithsonian Astrophysical Observatory Internal report, OAPA-R1-1994, 11 Gennaio, 1994, 1994.

[80] M. Barbera, “Soft X-ray/UV plastic film transmission measurements conducted at the Space Sciences Laboratory, University of California at Berkeley: data reduction, and results,” Smithsonian Astrophysical Observatory Internal report, OAPA-R2-1994, 10 Febbraio, 1994, 1994.

[81] M. Barbera, A. Collura, S. Sciortino & S. Serio, “Soft X-ray filter transmission measurements conducted at the X-ray Astronomy Calibration and Testing facility of the Osservatorio Astronomico di Palermo G.S.Vaiana.,” Osservatorio Astronomico di Palermo G.S.Vaiana, OAPA-R3-1994, 18 Febbraio 1994, 1994.

[82] M. Barbera, “Soft X-ray/UV transmission measurements iof Parylene-C plastic film,” Smithsonian Astrophysical Observatory Internal report, OAPA-R4-1994, 24 Maggio, 1994, 1994.

[83] M. Barbera, “Simulations of the HRC Ultraviolet count-rate due to bright UV sources,” Smithsonian Astrophysical Observatory Internal report, OAPA-R5-1994, 11 Giugno, 1994, 1994.

[84] M. Barbera & A. Collura, “Soft X-ray plastic film transmission measurements conducted at the XACT facility of the Osservatorio Astronomico di Palermo G.S.Vaiana.,” Osservatorio Astronomico di Palermo G.S.Vaiana, OAPA-R6-1994, 17 Giugno, 1994, 1994.

[85] E. Antonucci, M. A. Dodero, G. Peres, F. Reale, S. Serio & B.V. Somov, “Soft X-ray Line Shifts as Signature of the Flare Heating Process,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 159, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992.

[86] M. Barbera, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “Coronal Emission at Low-Mass End: Results from an Optical-Selected Sample of Nearby K and M Stars Surveyed with the Einstein Observatory ,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 279, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 .

[87] A. Ciaravella, G. Peres & S. Serio, “Thermal Transport in Static Stellar Coronal Loops,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 595, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992.

[88] A. Collura, M. Barbera, G. Inzerillo, A. Maggio, G. Micela, F. Mirabello, S. Sciortino & S. Serio, “The G.S. Vaiana X-ray Astronomy Calibration and Testing Facility (XACT),” Mem. S.A.It. 64 (1993), 633, Presented at the XXXVI Meeting of SAIt, Teramo 6-9 Maggio 1992.

[89] A. Collura, G. Micela, S. Sciortino, F. R. Harnden, Jr. & R. Rosner, “A Multi-Color Method for The Classification of Low Resolution X-ray Spectra,” Mem. S.A.It. 64 (1993), 637, Presented at the XXXVI Meeting of SAIt, Teramo 6-9 Maggio 1992.

[90] A. Collura, M. Barbera, G. Inzerillo, A. Maggio, G. Micela, F. Mirabello, S. Sciortino, S. Serio & G. Peres, “The G. S. Vaiana X-ray Astronomy Calibration and Testing (XACT) Facility,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 275, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992 .

[91] F. Damiani & G. Micela, “The X-ray Emission of T-Tauri Stars in Taurus-Auriga as Seen with the Einstein Observatory IPC: Properties and Relationship with PMS Stars Activity.,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 283, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26.

[92] F. Damiani, G. Micela & S. Sciortino, “X-ray Observations of Herbig Ae/Be Stars,” in The nature and evolutionary status of Herbig Ae/Be Stars, The, P. S., Perez, M. R. & van den Heuvel, E. P. J., eds., A.S.P. Conference Series (San Francisco, California), 1993, 291, Presented at the meeting “The nature and evolutionary status of Herbig Ae/Be Stars’, Amsterdam, 26-29 October 1993.

[93] F. Damiani, G. Micela & S. Sciortino, “Rosat PSPC Observations of HD200775, BD+61-154, and VV Ser.,” in The nature and evolutionary status of Herbig Ae/Be Stars, The, P. S., Perez, M. R. & van den Heuvel, E. P. J., eds., A.S.P. Conference Series (San Francisco, California), 1993, 301, Presented at the meeting “The nature and evolutionary status of Herbig Ae/Be Stars’, Amsterdam, 26-29 October 1993.

[94] F. Favata, G. Micela, S. Sciortino & G.S. Vaiana, “Coronal source log N(¿S) - log S modeling: a tool to study stellar X-ray emission.,” Mem. S.A.It. in press (1993), Presented at “Congresso Astronet 1990”, Cagliari Ottobre 4-5.

[95] F. Favata, G. Micela & S. Sciortino, “An Analysis of the Stellar Content of the Einstein Extended Medium Sensitivity Survey Using Age-Homogeneous X-ray Luminosity Functions.,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 287, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 .

[96] F. Favata, M. Barbera, G. Micela & S. Sciortino, “Lithium Abundance Determination in EMSS Stellar Sources,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 303, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 .

[97] A. Maggio, S. Sciortino, L. Bianchi, F. R. Harnden, Jr. & R. Rosner, “Evolution of Stellar Activity in Early Post-Main Sequence Phases,” in Inside the Stars, W. W. Weiss & A. Baglin, eds. #40, Conference Series ASP, 1993, 648, Presented at the IAU Colloquium 137 “Inside the Stars”, Vienna, April 13-17 1992.

[98] A. Maggio, S. Sciortino, L. Bianchi, F. R. Harnden, Jr. & R. Rosner, “The X-ray and UV emission from Post-Main Sequence Stars: The Connection between Surface activity and Evolution,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 291, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 .

[99] A. Maggio, S. Sciortino & F. R. Harnden, Jr., “Soft X-ray Emission and Spectra of Evolved Stars From ROSAT Pointed Observations,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 365, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 .

[100] S. Orlando, G. Peres & S. Serio, “Siphon Flow Models of Solar Coronal Loops,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 163, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992 .

[101] G. Peres, G. Micela, S. Sciortino & S. Serio, “A Tool for X-ray Sky Simulations,” Mem. S.A.It. 64 (1993), 385, Presented at the Workshop “Space Projects to Probe the Internal Structure and The Magnetic Activity of the Sun and Sun-like Stars”, Catania, 5-8 November 1991.

[102] F. Reale, S. Serio & G. Peres, “Scaling Laws For the Decay Phase of Stellar Loop Flares,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 599, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992.

[103] F. Reale, G. Peres & S. Serio, “Thermodynamic Decay of Stellar Coronal Flares,” Mem. SAIt. 64 (1993), 555, Presented at the XXXVI Meeting of SAIt, Teramo 6-9 Maggio 1992.

[104] F. Reale, M. Barbera, G. Peres, S. Sciortino & S. Serio, “Transputer Based Concurrent Calculations,” Mem. S.A.It. in press (1993), Presented at “Congresso Astronet 1990”, Cagliari Ottobre 4-5.

[105] A. Maggio, F. Reale & G. Peres, “The Palermo Analysis System for Astrophysical Plasmas (ASAP): Rationale, Development and Applications,” Mem. S.A.It. 64 (1993), 1023, Presented at the 3rd Astronet Meeting, Trieste.

[106] F. Reale, F. Bocchino & S. Sciortino, “Parallel Computations in Astrophysics: Using a Workstation Array,” Mem. S.A.It. 64 (1993), 1011, Presented at the 3rd Astronet Meeting, Trieste.

[107] S. Serio, F. Reale, G. Peres, J. Jakimiec, B. Sylwester & J. Sylwester, “Flare Evolution in the Density-Temperature Diagram,” 1993, Presented at the IAU colloquium 133, Iguazu, July 1991.

[108] B. Sylwester, J. Sylwester, M. Siarkowski, S. Serio & F. Reale, “Analysis of Flare Evolution in the Emission Measure-Temperature Diagrams For Selected Events Observed by SMM,” in Physics of Solar and Stellar Coronae, J. Linsky & S. Serio, eds., Kluwer Academic Publishers, 1993, 147, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26 1992.

[109] R. Ventura, I. Pagano, G. Peres & M. Rodono‘, Observations of Low Level Optical Variability in Active Stars, Kluwer Academic Publishers, 1993, Presented at the “G. S. Vaiana Memorial Symposium: Advances in Stellar and Solar Coronal Physics”, Palermo, June 22-26, 1992.

[110] F. Reale, G. Peres & S. Serio, “Idrodinamica di plasmi astrofisici,” in Scienza e supercalcolo al CINECA - Rapporto 1993, S. Bassini, G. Erbacci, G. Paruolo & M. Voli, eds., CINECA, 1993, 8, Rapporto 1993.

[111] R. Ventura, G. Peres, I. Pagano & M. Rodono‘, “Time Series Analysis of Low Amplitude Variability and Transients in Late Type Active Stars,” in Application of Time Series Analysis in Astronomy and Meteorology, 1993, in press, Proceedings of the International Conference ”Application of Time Series Analysis in Astronomy and Meteorology”, Padova, September 6-10, 1993.

[112] G. Peres, G. Micela, M. Barbera, S. Sciortino & S. Serio, “Sky simulations for X-ray Telescopes,” Mem. SAIt. 64 (1993), 716, Proceedings of the XXXVII Congress of the S.A.It - May 1993, Florence.

[113] G. Micela, S. Sciortino, F. R. Harnden, Jr., V. Kashyap & R. Rosner, “High Sensitivity ROSAT Observations of the Pleiades,” Mem. SAIt. 64 (1993), 695, Proceedings of the XXXVII Congress of the S.A.It - May 1993, Florence.

[114] F. Favata, G. Micela, S. Sciortino & M. Barbera, “The Correlation Between Lithium and Activity in Late-Type Main Sequence Stars,” in Cool Stars, Stellar Systems, and The Sun - Eight Cambridge Workshop, Caillault J. P., ed., A.S.P. Conference Series (San. Francisco, CA), 1993, 276, Presented at the Eighth Cool Star Workshop, Athene, Georgia, Oct. 11-14.

[115] F. Damiani, G. Micela, S. Sciortino & F. Favata, “X-Ray and H-alpha Observations of Rapidly Rotating Pre-main-Sequence Stars ,” in Cool Stars, Stellar Systems, and The Sun - Eight Cambridge Workshop, Caillault J. P., ed., A.S.P. Conference Series (San. Francisco, CA), 1993, 71, Presented at the Eighth Cool Star Workshop, Athene, Georgia, Oct. 11-14.

[116] G. Micela, S. Sciortino, V. Kashyap, F. R. Harnden, Jr. & R. Rosner, “An X-Ray Study of a Coeval Star Population: a Deep Rosat Observation of the Core of the Pleiades Cluster,” in Cool Stars, Stellar Systems, and The Sun - Eight Cambridge Workshop, Caillault J. P., ed., A.S.P. Conference Series (San. Francisco, CA), 1993, 110, Presented at the Eighth Cool Star Workshop, Athene, Georgia, Oct. 11-14.

[117] S. Sciortino, G. Micela, V. Kashyap, F. R. Harnden, Jr. & R. Rosner, “A Variability Study of X-Ray Emission of Pleiades Stars,” in Cool Stars, Stellar Systems, and The Sun - Eight Cambridge Workshop, Caillault J. P., ed., A.S.P. Conference Series (San. Francisco, CA), 1993, 140, Presented at the Eighth Cool Star Workshop, Athene, Georgia, Oct. 11-14.

[118] A. Maggio, S. Serio & G. Peres, “Stellar XUV Spectra from Coronal Loop Models,” Bullettin of AAS 25 (1993), 864, Proceedings of the 182nd AAS Meeting (Abstract Reports).

[119] C.-C. Cheng & R. Pallavicini, “X-ray Spectroscopic Diagnostics of the Hydrodynamics of Flares on M Dwarf Stars,” in UV and X-ray Spectroscopy of Astrophysical and Laboratory Plasmas, E.H. Silver & S.M. Kahn, eds., 1993, 525.

[120] R. Gratton, R. Bhatia, G. Bonanno, P. Bruno, A. Cali, S. Catalano, A. Cavazza, R. Claudi, R. Cosentino, N. Douglas, R. Merighi, P. Molaro, R. Pallavicini & R. Viotti, “A High Resolution Spectrograph For the Galileo National Telescope,” Memorie S.A.It. 63 (1993), 672.

[121No references found. ] .